# Deterministic factorization of constant-depth algebraic circuits in subexponential time

### v1

@@ 8.1 From irreducibility testing to divisibility tests
The following lemma is the first step towards our approach of relating irreducibility testing to divisibility tests.
@@
Let $P(T,\mathbf{x},z)$ be a nonzero squarefree polynomial that is monic in $z$ and $T$-regularized with respect to $z$ with $\deg_{z}P=D$. Let $k>\deg_{T}(P)$ and $P(0,\mathbf{x},z)$ (which equals $P(0,\mathbf{0},z)$) be squarefree.
Also, let $P(T,\mathbf{x},z)=\prod_{i\in[D]}(z-\varphi_{i}(T,\mathbf{x}))$ be the factorization of $P(T,\mathbf{x},z)$ where the roots $\varphi_{i}(T,\mathbf{x})$ are from the ring of power series $\mathbb{F}[\mathbf{x}]\llbracket T\rrbracket$. [^9] Such a factorization exists because each root of the univariate $P(0,\mathbf{x},z)$ can be extended to a power series root via Newton iteration; see [DSS22, Theorem 4]. Suppose $R_{1},R_{2},\ldots,R_{D}\in\mathbb{F}[T,\mathbf{x}]$ are truncated, non-degenerate, approximate $z$-roots of $P$ of order $k$ with respect to $T$ that satisfy
$R_{i}\equiv\varphi_{i}\mod T^{k}\,.$
For any subset $S\subseteq[D]$, define $Q_{S}(T,\mathbf{x},z):=(\prod_{i\in S}(z-R_{i}(T,\mathbf{x}))\;\mathrm{trunc}\;T^{k})$. Then, for any $S\subseteq[D]$, $Q_{S}(T,\mathbf{x},z)$ divides $P(T,\mathbf{x},z)$ if and only if $\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))\in\mathbb{F}[T,\mathbf{x},z]$ (and not just $\mathbb{F}[\mathbf{x}]\llbracket T\rrbracket[z]$). In other words, $Q_{S}$ divides $P$ if and only if $\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))$ is a polynomial in $T,\mathbf{x},z$.
Moreover, if $Q_{S}$ divides $P$ then, $Q_{S}$ equals the polynomial $\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))$.
@@
For any $S\subseteq[D]$, if $F_{S}(T,\mathbf{x},z):=\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))\in\mathbb{F}[T,\mathbf{x},z]$ i.e. $F_{S}(T,\mathbf{x},z)$ is a true polynomial factor of $P(T,\mathbf{x},z)$, then we observe that $Q_{S}(T,\mathbf{x},z)=(\prod_{i\in S}(z-R_{i}(T,\mathbf{x}))\bmod\left\langle T^{k}\right\rangle)$ must equal $F_{S}(T,\mathbf{x},z)$: since $k>\deg_{T}(P)\geq\deg_{T}(F_{S})$,
$F_{S}(T,\mathbf{x},z)$ $=F_{S}(T,\mathbf{x},z)\;\mathrm{trunc}\;\left\langle T^{k}\right\rangle$
$=\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))\;\mathrm{trunc}\;\left\langle T^{k}\right\rangle$
$=\prod_{i\in S}(z-R_{i}(T,\mathbf{x}))\;\mathrm{trunc}\;\left\langle T^{k}\right\rangle$
$=Q_{S}(T,\mathbf{x},z)\,.$
For the other direction, let $S\subseteq[D]$ and suppose $Q_{S}$ divides $P$. Thus, $Q_{S}=\prod_{i\in U}(z-\varphi_{i}(T,\mathbf{x}))$ for some $U\subseteq[D]$. In particular, $Q_{S}(0,{\bm{0}},z)=\prod_{i\in U}(z-\varphi_{i}(0,{\bm{0}}))=\prod_{i\in U}(z-R_{i}(0,{\bm{0}}))$, which follows because for each $i\in[D]$, $\varphi_{i}(T,\mathbf{x})=R_{i}(T,\mathbf{x})\bmod\left\langle T^{k}\right\rangle$ for some $k>\deg_{T}(P)\geq 1$. But, by definition of $Q_{S}$, $Q_{S}(0,{\bm{0}},z)=\prod_{i\in S}(z-R_{i}(0,{\bm{0}}))$ and so, $\prod_{i\in S}(z-R_{i}(0,{\bm{0}}))=\prod_{i\in U}(z-R_{i}(0,{\bm{0}}))$. Since $P(0,{\bm{0}},z)$ is squarefree [^10] While squarefreeness implies that each irreducible factor of $P(0,{\bm{0}},z)$ over $\mathbb{F}$ is unique, this also implies that each root of $P(0,{\bm{0}},z)$ in $\overline{\mathbb{F}}$ is unique because squarefreeness is captured by the non-vanishing of the discriminant, which only depends on the coefficients of $P(0,{\bm{0}},z)$ and not on the underlying field., each $\varphi_{i}(0,{\bm{0}})$ is distinct, and equivalently, each $R_{i}(0,{\bm{0}})$ is distinct. Thus, $U=S$ and $\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))\in\mathbb{F}[T,\mathbf{x},z]$.
An application of the above lemma that we use later in this section is the lemma below. This lemma characterizes the irreducible factors of a polynomial $P$ by a set of (potentially exponentially many) divisibility tests. We later reduce these divisibility tests into instances that we derandomize to get Theorem 5.1.
@@
Let $P(T,\mathbf{x},z)$ be a nonzero squarefree polynomial that is monic in $z$ and $T$-regularized, with $\deg_{z}P=D$. Let $k>\deg_{T}(P)$ and $P(0,\mathbf{x},z)$ (which equals $P(0,\mathbf{0},z)$) be squarefree.
Also, let $P(T,\mathbf{x},z)=\prod_{i\in[D]}(z-\varphi_{i}(T,\mathbf{x}))$ be the factorization of $P(T,\mathbf{x},z)$ where the roots $\varphi_{i}(T,\mathbf{x})$ are from the ring of power series $\mathbb{F}[\mathbf{x}]\llbracket T\rrbracket$. [^11] Such a factorization exists because each root of the univariate $P(0,\mathbf{x},z)$ can be extended to a power series root via Newton iteration; see [DSS22, Theorem 4]. Suppose $R_{1},R_{2},\ldots,R_{D}\in\mathbb{F}[T,\mathbf{x}]$ are truncated, non-degenerate, approximate $z$-roots of $P$ of order $k$ that satisfy
$R_{i}\equiv\varphi_{i}\mod T^{k}\,.$
Then, for any subset $S\subseteq[D]$, $Q_{S}(T,\mathbf{x},z):=\prod_{i\in S}(z-R_{i}(T,\mathbf{x}))\;\mathrm{trunc}\;T^{k}$ is an irreducible factor of $P(T,\mathbf{x},z)$ if and only if
  • $Q_{S}(T,\mathbf{x},z)$ divides $P(T,\mathbf{x},z)$, and
  • for every strict subset $U\subset S$ with $|U|\geq 1$, $Q_{U}(T,\mathbf{x},z)\nmid P(T,\mathbf{x},z)$, where $Q_{U}(T,\mathbf{x},z):=\prod_{i\in U}(z-R_{i}(T,\mathbf{x}))\;\mathrm{trunc}\;T^{k}$
@@
For the forward direction of the proof, suppose $Q_{S}$ is an irreducible factor of $P$. Then clearly, $Q_{S}$ must divide $P$. Moreover, from 8.1, we get that $Q_{S}$ equals $\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))$. To obtain the second item, we argue via contradiction. Suppose that there exists a proper subset $U$ of $S$ such that $Q_{U}$ divides $P$. So, from 8.1, we have that $\prod_{i\in U}(z-\varphi_{i}(T,\mathbf{x}))$ must be a polynomial in $T,\mathbf{x},z$ and equal $Q_{U}$. But then, from the same lemma (invoked with the polynomial $P$ being replaced by $Q_{S}$), we have that $Q_{U}$ must be a factor of $Q_{S}$. Since $U$ is a non-empty proper subset of $S$, the $z$-degree of $Q_{U}$ is at least one and strictly less than the $z$-degree of $Q_{S}$, and hence $Q_{U}$ is a non-trivial factor of $Q_{S}$. But this contradicts the irreducibility of $Q_{S}$. This completes the proof of the second item.
For the reverse direction, the first item in the hypothesis already guarantees that $Q_{S}$ is a factor of $P$; Lemma 8.1 further tells us that $Q_{S}=\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))$. Now, all we need to argue is the irreducibility of $Q_{S}$. We do this via contradiction. Suppose that $Q_{S}$ is not irreducible, and let $A$ be a (non-trivial) factor of $Q_{S}$. Since $Q_{S}$ is monic in $z$, we get that $A$ is also monic in $z$ with $z$-degree at least one and strictly less than $|S|$. So, by an application of 8.1, we get that there is a non-trivial proper subset $U$ of $S$ such that $A=\prod_{i\in U}(z-\varphi_{i}(T,\mathbf{x}))$ and moreover, $A=\prod_{i\in U}(z-R_{i}(T,\mathbf{x}))\;\mathrm{trunc}\;T^{k}$. But since $A$ divides $Q_{S}$ and $Q_{S}$ divides $P$, we get that $A$ divides $P$, which immediately contradicts the second item in the hypothesis.
This completes the proof of the lemma. ∎
An immediate consequence of the above lemma is the following corollary that gives a certificate of irreducibility of a polynomial, in terms of an exponentially large set of divisibility tests.
@@
Let $P(T,\mathbf{x},z)$ be a nonzero squarefree polynomial that is monic in $z$ and $T$-regularized, with $\deg_{z}P=D$. Let $k>\deg_{T}(P)$ and $P(0,\mathbf{x},z)$ (which equals $P(0,\mathbf{0},z)$) be squarefree.
Let $P(T,\mathbf{x},z)=\prod_{i\in[D]}(z-\varphi_{i}(T,\mathbf{x}))$ be the factorization of $P(T,\mathbf{x},z)$ where the roots $\varphi_{i}(T,\mathbf{x})$ are from the ring of power series $\mathbb{F}[\mathbf{x}]\llbracket T\rrbracket$.
Then, $P$ is irreducible over $\mathbb{F}$ if and only if for every subset $S$ of $[D]$, the polynomial $Q_{S}(T,\mathbf{x},z):=\prod_{i\in S}(z-R_{i}(T,\mathbf{x}))\;\mathrm{trunc}\;T^{k}$ does not divide $P$ in the ring $\mathbb{F}[T,\mathbf{x},z]$.
The above corollary together with the reduction from divisibility testing to PIT in the next section gives us a reduction from irreducibility testing to PIT. However, we note that 8.3 gives exponentially many (in degree $D$) divisibility testing instances, and thus it isn’t immediately clear if this reduction can be useful for obtaining subexponential time irreducible testing algorithms. It is for this reason that our algorithm for factorization does not proceed directly using this reduction and only uses the corollary (and the previous lemmas) in its analysis.
@@ 8.2 From divisibility testing to polynomial identity tests
This almost immediately follows from 8.6 since each of the polynomials
@@
From 8.5, we get that there is a multi-output constant-depth circuit $C$ on $\ell$ variables that can be computed by a constant-depth circuit of size and degree $\operatorname{poly}(\ell)$ such that
@@
The following is an immediate consequence of 8.8 and 8.7.
@@
We take the circuits for $h_{i}(T,\mathbf{w})$ — given by 8.11, denoted by $V_{i}(T,\mathbf{w})$ — and consider the circuit $Q(V_{0}(\kappa,\mathbf{w}),V_{1}(\kappa,\mathbf{w}),\dots,V_{d}(\kappa,\mathbf{w}))$. Clearly, it is of depth $O(\Delta)$ and size $s^{\prime}\leq\operatorname{poly}(s,\deg(C^{\prime}),k)\cdot(\rho\log k)^{\operatorname{poly}(d)}$ and satisfies
@@
Similar to 7.1121212Refer to the part of the proof where we show that $h_{i}(T,\mathbf{w})$ can be computed by constant-depth circuits., the claim now follows from the definition of $B$, Corollary 3.4 and the fact that the degree of $B$ in $T$ is at most $\deg(C^{\prime})$. ∎
@@
Suppose $P(T,\mathbf{x},z)=\prod_{i=1}^{D}(z-\varphi_{i}(T,\mathbf{x}))$, where each $\varphi_{i}(T,\mathbf{x})$ is a power series root from $\mathbb{F}[\mathbf{x}]\llbracket T\rrbracket$. Then, by 8.1 and 8.2, we have that for any subset $S\subseteq[D]$, $F_{S}(T,\mathbf{x},z):=\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))$ is an irreducible polynomial factor of $P(T,\mathbf{x},z)$ if and only if
@@
  • for every strict subset $U\subset S$, $Q_{U}(T,\mathbf{x},z)\nmid P(T,\mathbf{x},z)$
@@
  • for every strict subset $U\subset S$, $Q_{U}(T,\mathbf{KI}_{g_{\sigma},\mathcal{S}}(\mathbf{w}),z)\nmid P(T,\mathbf{KI}_{g_{\sigma},\mathcal{S}}(\mathbf{w}),z)$
6.1 tells us that $P(T,\mathbf{KI}_{g_{\sigma},\mathcal{S}}(\mathbf{w}),z)$ will remain a nonzero squarefree [^14] Squarefreeness is equivalent to each root being non-degenerate, which is maintained by 6.1. polynomial that is monic in $z$ and $T$-regularized, with $\deg_{z}P=D$. Furthermore, $P(0,\mathbf{KI}_{g_{\sigma},\mathcal{S}}(\mathbf{w}),z)=P(0,\mathbf{0},z)$ will be squarefree. Thus, we can apply 8.1 and 8.2 to the above two conditions to get that for any subset $S\subseteq[D]$, $F_{S}(T,\mathbf{x},z)=\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))$ is an irreducible polynomial factor of $P(T,\mathbf{x},z)$ if and only if
@@
  • for every strict subset $U\subset S$, $\prod_{i\in T}(z-\varphi_{i}(T,\mathbf{KI}_{g_{\sigma},\mathcal{S}}(\mathbf{w})))\not\in\mathbb{Q}[T,\mathbf{w},z]$

### v2

@@ 8.1 From irreducibility testing to divisibility tests
Throughout this section, we assume that $P(T,\mathbf{x},z)$ is a nonzero squarefree polynomial that is monic in $z$, $T$-regularised with respect to $z$, and it satisfies that $P(0,\mathbf{x},z)=P(0,{\bm{0}},z)$ is squarefree. Let $D=\deg_{z}P$ and let $P(T,\mathbf{x},z)=\prod_{i\in[D]}(z-\varphi_{i}(T,\mathbf{x}))$ be the factorization [^9] Such a factorization exists because each root of the univariate $P(0,\mathbf{x},z)$ can be extended to a power series root via Newton iteration; see [DSS22, Theorem 4]. of $P(T,\mathbf{x},z)$ where $\varphi_{i}(T,\mathbf{x})$ are from the ring of power series $\overline{\mathbb{F}}[\mathbf{x}]\llbracket T\rrbracket$.
@@
Let $P(T,\mathbf{x},z)$ and $\varphi_{1}(T,\mathbf{x}),\ldots,\varphi_{D}(T,\mathbf{x})$ be defined as above, and let $k>\deg_{T}(P)$. For any $G(T,\mathbf{x},y)\in\overline{\mathbb{F}}(T,\mathbf{x},z)$ that divides $P(T,\mathbf{x},z)$, there is some subset $U\subseteq[D]$ such that
$G(T,\mathbf{x},y)=\prod_{i\in U}(z-\varphi_{i}(T,\mathbf{x}))\;\mathrm{trunc}\;T^{k}.$
and $\prod_{i\in U}(z-\varphi_{i}(T,\mathbf{x}))$ is in fact a polynomial (and not just a power series) that equals $G(T,\mathbf{x},y)$.
@@
Since $P(T,\mathbf{x},z)$ is squarefree and we have the factorisation $P(T,\mathbf{x},z)=\prod_{i\in[D]}(z-\varphi_{i}(T,\mathbf{x}))$ in $\mathbb{F}[\mathbf{x},z]\llbracket T\rrbracket$ and hence any polynomial factor $G(T,\mathbf{x},z)$ must be $\prod_{i\in U}(z-\varphi_{i}(T,\mathbf{x}))$ for some $U\subseteq[D]$. Since $\deg_{T}G\leq\deg_{T}P<k$, we have
$G(T,\mathbf{x},y)$ $=G\;\mathrm{trunc}\;T^{k}$
$=\prod_{i\in U}(z-\varphi_{i}(T,\mathbf{x}))\;\mathrm{trunc}\;T^{k}\qed$
For our algorithm, we would be interested in divisibility over $\mathbb{F}$ and not over $\overline{\mathbb{F}}$. The following lemma gives some sufficient conditions for when the above polynomial $Q_{S}(T,\mathbf{x},z)$ has coefficients in $\mathbb{F}$.
@@
Let $P(T,\mathbf{x},z)$ be a polynomial and $\varphi_{1}(T,\mathbf{x}),\ldots,\varphi_{D}(T,\mathbf{x})$ be power series as defined above. For each $i\in[D]$, let $\alpha_{i}=\varphi_{i}(0,{\bm{0}})$, and let $\mathcal{Z}=\left\{\alpha_{1},\ldots,\alpha_{D}\right\}$ (which are the roots of $P(0,\mathbf{x},z)=P(0,{\bm{0}},z)$ in the algebraic closure $\overline{\mathbb{F}}$). Fix some $k\geq 0$. For a subset $S\subset[D]$, let
$Q_{S}(T,\mathbf{x},z):=\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))\;\mathrm{trunc}\;T^{k}$
for some $k\geq 0$.
If the polynomial $g_{S}(z)=\prod_{i\in S}(z-\alpha_{i})=Q_{S}(0,{\bm{0}},z)=Q_{S}(0,\mathbf{x},z)$ has all its coefficients in the base field $\mathbb{F}$, then so does $Q_{S}(T,\mathbf{x},z)$.
@@
Define the following rational functions $\Gamma^{(i)}(T,\mathbf{x},\zeta)$ for $i\geq 0$ as follows:
$\Gamma^{(0)}(\zeta)$ $:=\zeta$
$\Gamma^{(i+1)}(\zeta)$ $:=\Gamma^{(i)}(\zeta)-\frac{P(T,\mathbf{x},\Gamma^{(i)}(\zeta))}{\partial_{z}P(0,{\bm{0}},\zeta)}\quad\text{for $i>0$}.$
Note that $\Gamma^{(i)}(\alpha)$ is in fact a polynomial in $T,\mathbf{x}$ for every $\alpha$ such that $\partial_{z}P(0,{\bm{0}},\alpha)\neq 0$. By Lemma 4.5, for any $\alpha_{i}\in\mathcal{Z}$, we have that $\Gamma^{(k)}(\alpha_{i})=\varphi_{i}\bmod{T^{k}}$ and therefore
$Q_{S}(T,\mathbf{x},z)=\prod_{i\in S}(z-\Gamma^{(k)}(\alpha_{i}))\;\mathrm{trunc}\;T^{k}$
Therefore, every coefficient of $Q_{S}(T,\mathbf{x},z)$ is a symmetric function of the set $\left\{\alpha_{i}\ :\ i\in S\right\}$. Since $g(z)=\prod_{i\in S}(z-\alpha_{i})$ is assumed to have coefficients in the base field $\mathbb{F}$, this implies that every elementary symmetric function of the set $\left\{\alpha_{i}\ :\ i\in S\right\}$ is in $\mathbb{F}$. By the fundamental theorem of symmetric polynomials, this implies that every coefficient of $Q_{S}(T,\mathbf{x},z)$ is also in $\mathbb{F}$. ∎
With the above two lemmas, we can now characterise the factors (over $\mathbb{F}$) of $P(T,\mathbf{x},z)$ using the polynomials $Q_{S}(T,\mathbf{x},z)$ defined above.
Lemma 8.3.
Let $P(T,\mathbf{x},z)$ and $\varphi_{1}(T,\mathbf{x}),\ldots,\varphi_{D}(T,\mathbf{x})$ be defined as above, and let $k>\deg_{T}P$. Let $\alpha_{i}=\varphi_{i}(0,{\bm{0}})$ for each $i\in[D]$. For a subset $S\subseteq[D]$, define the polynomial
$Q_{S}(T,\mathbf{x},z)=\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))\;\mathrm{trunc}\;T^{k}.$
Then, the set of factors of $P$ over the base field $\mathbb{F}$ is the same as
$\left\{Q_{U}(T,\mathbf{x},z)\in\overline{\mathbb{F}}[T,\mathbf{x},z]\ :\ \begin{array}[]{c}U\subseteq[D]\text{ with }Q_{U}(T,\mathbf{x},z)\text{ dividing }P\\ \text{and }Q_{U}(0,\mathbf{x},z)=Q_{U}(0,{\bm{0}},z)\in\mathbb{F}[z]\end{array}\right\}$
Proof.
Any $G(T,\mathbf{x},z)\in\mathbb{F}[T,\mathbf{x},z]$ that is a factor of $P(T,\mathbf{x},z)$, by Lemma 8.1, is equal to $Q_{U}(T,\mathbf{x},z)=\prod_{i\in U}(z-\varphi_{i}(T,\mathbf{x}))$ for some $U\subseteq[D]$. Additionally, if $G(T,\mathbf{x},z)\in\mathbb{F}[T,\mathbf{x},z]$, then $G(0,\mathbf{x},z)=Q_{U}(0,\mathbf{x},z)\in\mathbb{F}[z]$.
For the other direction, consider any $Q_{U}(T,\mathbf{x},z)\in\overline{\mathbb{F}}[T,\mathbf{x},z]$ that divides $P(T,\mathbf{x},z)$ and also satisfies that $Q_{U}(0,{\bm{0}},z)=\prod_{i\in U}(z-\alpha_{i})$ is a polynomial with coefficients in $\mathbb{F}$. By Lemma 8.2, we have that $Q_{U}(T,\mathbf{x},z)$ must also be a polynomial in $\mathbb{F}[T,\mathbf{x},z]$ and hence is a factor of $P(T,\mathbf{x},z)$ over $\mathbb{F}$. ∎
This immediately implies the following certificate for irreducibility.
@@
Let $P(T,\mathbf{x},z)\in\mathbb{F}[T,\mathbf{x},z]$ and $\varphi_{1}(T,\mathbf{x}),\ldots,\varphi_{D}(T,\mathbf{x})$ be defined as above, and let $k>\deg_{T}(P)$. Then $P(T,\mathbf{x},z)$ is irreducible over $\mathbb{F}$ if and only if for every $\emptyset\neq S\subsetneq[D]$ such that $\prod_{i\in S}(z-\alpha_{i})$ has all coefficients in $\mathbb{F}$, we have that $Q_{S}(T,\mathbf{x},z)\in\mathbb{F}[T,\mathbf{x},z]$ does not divide $P(T,\mathbf{x},z)$ where
$Q_{S}(T,\mathbf{x},z)=\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))\;\mathrm{trunc}\;T^{k}.$
Remark 8.5.
Essentially the same proof as Lemma 8.3 also characterizes the set of factors of $P(T,\mathbf{x},z)$ over the algebraic closure $\overline{\mathbb{F}}$, and thus gives a certificate for absolute irreducibility using divisibility tests over $\overline{\mathbb{F}}$.
Corollary 8.6.
Let $P(T,\mathbf{x},z)\in\mathbb{F}[T,\mathbf{x},z]$ and $\varphi_{1}(T,\mathbf{x}),\ldots,\varphi_{D}(T,\mathbf{x})$ be defined as above, and let $k>\deg_{T}(P)$. Then $P(T,\mathbf{x},z)$ is absolutely irreducible if and only if for every $\emptyset\neq S\subsetneq[D]$, we have that $Q_{S}(T,\mathbf{x},z)\in\overline{\mathbb{F}}[T,\mathbf{x},z]$ does not divide $P(T,\mathbf{x},z)$ where
$Q_{S}(T,\mathbf{x},z)=\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))\;\mathrm{trunc}\;T^{k}.$
8.6 together with the reduction from divisibility testing to PIT in the next section gives us a reduction from irreducibility testing to PIT. However, we note that gives exponentially many (in degree $D$) divisibility testing instances, and thus it isn’t immediately clear if this reduction can be useful for obtaining subexponential time irreducible testing algorithms. It is for this reason that our algorithm for factorization does not proceed directly using this reduction and only uses the corollary (and the previous lemmas) in its analysis.
@@ 8.2 From divisibility testing to polynomial identity tests
This almost immediately follows from 8.9 since each of the polynomials
@@
From 8.8, we get that there is a multi-output constant-depth circuit $C$ on $\ell$ variables that can be computed by a constant-depth circuit of size and degree $\operatorname{poly}(\ell)$ such that
@@
The following is an immediate consequence of 8.11 and 8.10.
@@
We take the circuits for $h_{i}(T,\mathbf{w})$ — given by 8.14, denoted by $V_{i}(T,\mathbf{w})$ — and consider the circuit $Q(V_{0}(\kappa,\mathbf{w}),V_{1}(\kappa,\mathbf{w}),\dots,V_{d}(\kappa,\mathbf{w}))$. Clearly, it is of depth $O(\Delta)$ and size $s^{\prime}\leq\operatorname{poly}(s,\deg(C^{\prime}),k)\cdot(\rho\log k)^{\operatorname{poly}(d)}$ and satisfies
@@
Similar to 7.1101010Refer to the part of the proof where we show that $h_{i}(T,\mathbf{w})$ can be computed by constant-depth circuits., the claim now follows from the definition of $B$, Corollary 3.4 and the fact that the degree of $B$ in $T$ is at most $\deg(C^{\prime})$. ∎
@@
Suppose $P(T,\mathbf{x},z)=\prod_{i=1}^{D}(z-\varphi_{i}(T,\mathbf{x}))$, where each $\varphi_{i}(T,\mathbf{x})$ is a power series root from $\mathbb{F}[\mathbf{x}]\llbracket T\rrbracket$. Then, by 8.6, we have that for any subset $S\subseteq[D]$, $F_{S}(T,\mathbf{x},z):=\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))$ is an irreducible factor of $P(T,\mathbf{x},z)$ in $\mathbb{F}[T,\mathbf{x},z]$ if and only if
@@
  • for every $\emptyset\neq U\subsetneq S$ such that $\prod_{i\in U}(z-\alpha_{i})$ has all coefficients in $\mathbb{F}$, we have that $Q_{U}(T,\mathbf{x},z)\in\mathbb{F}[T,\mathbf{x},z]$ does not divide $P(T,\mathbf{x},z)$
@@
  • for every $\emptyset\neq U\subsetneq S$ such that $\prod_{i\in U}(z-\alpha_{i})$ has all coefficients in $\mathbb{F}$, we have that $Q_{U}(T,\mathbf{KI}_{g_{\sigma},\mathcal{S}}(\mathbf{w}),z)$ does not divide $P(T,\mathbf{KI}_{g_{\sigma},\mathcal{S}}(\mathbf{w}),z)$
6.1 tells us that $P(T,\mathbf{KI}_{g_{\sigma},\mathcal{S}}(\mathbf{w}),z)$ will remain a nonzero squarefree [^12] Squarefreeness is equivalent to each root being non-degenerate, which is maintained by 6.1. polynomial that is monic in $z$ and $T$-regularized, with $\deg_{z}P=D$. Furthermore, $P(0,\mathbf{KI}_{g_{\sigma},\mathcal{S}}(\mathbf{w}),z)=P(0,\mathbf{0},z)$ will be squarefree. Thus, we can apply 8.6 to the above two conditions to get that for any subset $S\subseteq[D]$, $F_{S}(T,\mathbf{x},z)=\prod_{i\in S}(z-\varphi_{i}(T,\mathbf{x}))$ is an irreducible polynomial factor of $P(T,\mathbf{x},z)$ if and only if
@@
  • for every $\emptyset\neq U\subsetneq S$ such that $\prod_{i\in U}(z-\alpha_{i})$ has all coefficients in $\mathbb{F}$, we have that $\prod_{i\in T}(z-\varphi_{i}(T,\mathbf{KI}_{g_{\sigma},\mathcal{S}}(\mathbf{w})))\not\in\mathbb{Q}[T,\mathbf{w},z]$